OpenStudy (anonymous):

[1.05] Use the distributive property to simplify the expression below: -6(2u - 1)

OpenStudy (anonymous):

Distributive property is: a(b+c) = ab + ac Can you apply that to solve this question?

OpenStudy (kainui):

The distributive property says you multiply something by every term in parenthesis. So following what anemonix wrote, here's an example: 2(3+4)=6+8=14 we can test it ourselves to make sure 2(3+4)=2(7)=14